OpenStudy (anonymous):

sin^2 θ (1 + cot^2 θ)

OpenStudy (solomonzelman):

\[\sin^2 θ (1 + \cot^2 θ) \] we know that cot^2x+1=csc^2x, So, \[\sin^2 θ(\csc^2 θ)\] sin is inverse of csc so it would be just 1

OpenStudy (anonymous):

thank you